# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v1

@@ C.2 Comparison against TAM
Table 7: Results for CVRP instances used in TAM. The result for AM and TAM is drawn from Hou et al. (2022), and the result for DeepACO is drawn from Ye et al. (2023).
@@
Appendix D Diversity Calculation
The diversity of $K$ solutions generated from a solver is defined as the average pairwise Jaccard distance, $\frac{1}{K(K-1)}\sum_{i}\sum_{j}d_{J}(x_{i},x_{j})$, where $d_{J}$ is defined as follows:
where $E_{x}$ is the set of edges composing the solution $x$. In TSP or CVRP, the (undirected) edge is a set of adjoining nodes in the solution path.
@@ F.1 Deep Constructive Policy for Neural Combinatorial Optimization
Constructive policy methods systematically build solutions by incrementally adding elements, beginning with an empty set and culminating in a complete solution. This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing (Nazari et al., 2018), graph combinatorial optimization (Khalil et al., 2017), and scheduling (Zhang et al., 2020).
In deep learning, various techniques adopt this constructive policy framework. Notably, the attention model (AM) (Kool et al., 2018) has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training. Subsequent advancements in AM have included enhanced decoder architecture (Son et al., 2023a; Xin et al., 2021; Luo et al., 2023), improved REINFORCE (Williams, 1992) baseline assessments (Kwon et al., 2020; Kim et al., 2022b), leveraging symmetries for sample efficiency (Kim et al., 2023a), the application of population-based reinforcement learning (RL) (Grinsztajn et al., 2022), and refined solution-building methods (Drakulic et al., 2023). Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally (Kim et al., 2021b; Hou et al., 2022; Ye et al., 2024). Additionally, AM-driven constructive policies have been effectively applied in hardware routing (Kim et al., 2020, 2021a) and placement tasks (Park et al., 2022; Kim et al., 2023b).
Graph combinatorial optimization, particularly in problems like maximum independent sets, has diversified from contrive VRPs policies due to their local decomposability and the relative ease of training value functions for partial solutions where VRPs are not locally decomposable due to their global contains. Starting with the influential works by Khalil et al. (2017), which employed deep Q networks (Mnih et al., 2013) for graph policy training, creating solutions constructively, Ahn et al. (2020) introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems. Zhang et al. (2023a) further developed this concept, training their MDP with forward-looking GFlowNets (Pan et al., 2023a).
@@ F.2 Deep Improvement Policy for Neural Combinatorial Optimization
The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt (Croes, 1958). Its advantage lies in achieving comparable performances to constructive policies, especially when allotted sufficient time for extensive improvement iterations. Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning (d O Costa et al., 2020; Hottung & Tierney, 2019; Ma et al., 2021, 2022, 2023; Kim et al., 2022a). Also, recent works have explored active search to refine solutions via policy adjustments (Hottung et al., 2021; Son et al., 2023b). Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes (Sun et al., 2023). This approach serves as a complementary extension to deep constructive policies, as it often enhances solutions initially formulated by constructive methods. Additionally, our method, which integrates local search into the training loop, could see further enhancements if advancements in this area are incorporated into our training procedures.
@@ F.3 Heatmap-based Neural Combinatorial Optimization
The heatmap-based method (Joshi et al., 2021) entails the acquisition of a graph representation, referred to as the “Heatmap,” from input graph instances. This Heatmap serves as a prior distribution, influencing the generation of a solution. However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search (Joshi et al., 2021), dynamic programming (Kool et al., 2022), or Monte Carlo tree search (Fu et al., 2021; Qiu et al., 2022; Sun & Yang, 2023). The key advantage of the heatmap-based method lies in its exceptional generalizability when compared to constructive policies. It operates at an abstract level, free from the constraints inherent in generating actual combinatorial solutions.

### v2

@@ D.5 Comparison against TAM
Table 9: Results for CVRP instances used in TAM. The results for AM and TAM is drawn from Hou et al. [24], and the result for DeepACO is drawn from Ye et al. [79].
@@
We compare GFACS and DeepACO against CVRP instances used in TAM [24]. These instances have a different capacity constraint from those we used for the main experiment (Table 1). Please refer to [24] for more details regarding the instance generation. Moreover, to match the processing time to be comparative to DeepACO, we set $K$ (equal to the number of ants for ACO) to 16.
D.6 Ablation study
Table 10: Results of the ablation study. The mean and standard deviation are reported based on three independent runs. The reported gap is calculated against Concorde [2] for TSP and PyVRP [77] for CVRP.
TSP500 CVRP500 BPP120
Method Obj. Gap (%) Obj. Gap (%) Obj.
GFACS 16.0842 $\pm$ 0.0043 1.56 64.1858 $\pm$ 0.0071 1.95 51.8428 $\pm$ 0.0706
   – Off-policy Training 17.0861 $\pm$ 0.0028 3.26 64.4248 $\pm$ 0.0155 2.33 52.3158 $\pm$ 0.1171
   – Energy Reshaping 16.8718 $\pm$ 0.0049 1.97 64.4121 $\pm$ 0.0069 2.31 52.0654 $\pm$ 0.0998
   – Shared $\varepsilon$-norm 17.5209 $\pm$ 0.0062 5.87 65.5069 $\pm$ 0.0171 4.05 90.7332 $\pm$ 9.1649
    + VarGrad [82] 16.8570 $\pm$ 0.0086 1.88 64.2771 $\pm$ 0.0293 2.09 51.9504 $\pm$ 0.1222
In this section, we present the comprehensive results of our ablation study, incorporating standard deviations from multiple independent runs (three runs) to ensure statistical significance. The findings demonstrate that each component of our technical proposals contributes to improvement, supported by empirical and statistically rigorous evidence.
@@ F.1 Deep constructive policy for neural combinatorial optimization
Constructive policy methods systematically build solutions by incrementally adding elements, beginning with an empty set and culminating in a complete solution. This approach ensures compliance with constraints by progressively narrowing the scope of actions, making it versatile for application in domains like vehicle routing [57, 41], graph combinatorial optimization [30, 1, 84], and scheduling [81].
In deep learning, various techniques adopt this constructive policy framework. Notably, the attention model (AM) [41] has been pivotal in vehicle routing problems (VRPs), establishing a new standard for stabilizing REINFORCE training. Subsequent advancements in AM have included enhanced decoder architecture [67, 78, 47], improved REINFORCE [76] baseline assessments [43, 37], leveraging symmetries for sample efficiency [32], the application of population-based reinforcement learning (RL) [21], ensemble-based learning [27, 86] and refined solution-building methods [16]. Recent advancements in hierarchical algorithms exploit AM-based policies for solution optimization: one policy broadly generates solutions while another refines them locally [36, 24, 80]. Additionally, AM-driven constructive policies have been effectively applied in hardware routing [34, 35] and placement tasks [60, 31].
Graph combinatorial optimization, particularly in problems like maximum independent sets, has diversified from contrive VRPs policies due to their local decomposability and the relative ease of training value functions for partial solutions where VRPs are not locally decomposable due to their global contains. Starting with the influential works by Khalil et al. [30], which employed deep Q networks [55] for graph policy training, creating solutions constructively, Ahn et al. [1] introduced an innovative Markov Decision Process (MDP) to generate multiple solution components per iteration by exploiting the local decomposability of graph combinatorial problems. Zhang et al. [84] further developed this concept, training their MDP with forward-looking GFlowNets [59].
@@ F.2 Deep improvement policy for neural combinatorial optimization
The improvement policy focuses on refining complete solutions, similar to local search methods like 2opt [11]. Its advantage lies in achieving comparable performances to constructive policies, especially when allotted sufficient time for extensive improvement iterations. Researchers typically adapt existing heuristic local search techniques, creating learning policies that emulate these methods through maximizing expected returns via reinforcement learning [12, 22, 48, 49, 50, 33]. Also, recent works have explored active search to refine solutions via policy adjustments [23, 66] or sampling the conditions from a continuous latent space [7]. Moreover, alternative strategies using discrete Langevin sampling instead of neural networks have shown promising outcomes [69]. This approach serves as a complementary extension to deep constructive policies, as it often enhances solutions initially formulated by constructive methods. Additionally, our method, which integrates local search into the training loop, could see further enhancements if advancements in this area are incorporated into our training procedures.
@@ F.3 Heatmap-based neural combinatorial optimization
The heatmap-based method [29] entails the acquisition of a graph representation, referred to as the “Heatmap," from input graph instances. This Heatmap serves as a prior distribution, influencing the generation of a solution. However, this approach is not self-contained and requires an additional solution construction step, which may involve techniques like beam search [29], dynamic programming [42], Monte Carlo tree search [20, 62, 70] or greedy search with a guide using an objective function [44]. The key advantage of the heatmap-based method lies in its exceptional generalizability when compared to constructive policies. It operates at an abstract level, free from the constraints inherent in generating actual combinatorial solutions.
@@
Broader Impact
We proposed a new probabilistic approach that can be trained using off-policy learning and applied to various combinatorial optimization problems, such as vehicle routing and scheduling. This method is relevant to industrial engineering, particularly in optimizing transportation and manufacturing processes, and can potentially extend to scientific discovery through chemical structure and biological sequence optimization. It is important to consider the ethical implications of probabilistic optimization methods, as they carry the potential for misuse in fields such as harmful drug discovery. While these algorithms offer several benefits, their application should be carefully monitored to ensure responsible and ethical use, minimizing the risk of unintended negative consequences.